OpenStudy (anonymous):

4 times the measure of the complement of an angle is equal to 2/3 the measure of the original angle's supplement. What is the measure of the angle?

OpenStudy (anonymous):

18° 48° 72° 108°

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@Keigh2015

OpenStudy (keigh2015):

Let the angle be x. Complement of x = 90 - x Supplement of x = 180 - x 4(90 - x) = 180 - x Find x from this...

OpenStudy (anonymous):

so i distribute it

OpenStudy (keigh2015):

yeah but tell me what you get before you turn it in

OpenStudy (anonymous):

so it would B

OpenStudy (keigh2015):

No

OpenStudy (anonymous):

how would i solve for X

OpenStudy (keigh2015):

let the angle be x. then complement of that = 90-x so, 4 times of complement = 4(90-x) 2/3 of original angle's supplement= 2/3 (180-x) so, 4(90-x)= 2/3 (180-x) cross multiply: 12(90-x) = 360 - 2x 12*90 - 12x = 360 - 2x 1080- 360= 12x-2x

OpenStudy (keigh2015):

Now you have to figure out the last part to get your answer

OpenStudy (anonymous):

C!

OpenStudy (anonymous):

thank yyou !

OpenStudy (keigh2015):

Correct

OpenStudy (keigh2015):

Your welcome

OpenStudy (mathstudent55):

angle is x complement is 90 - x supplement is 180 - x Now we follow the directions: 4 times the measure of the complement means 4(90 - x) that is equal to 2/3 of the supplement which means (2/3)(180 - x) \(4(90 - x) = \dfrac{2}{3}(180 - x) \)

OpenStudy (anonymous):

ty @mathstudent55 but i got it ! thank you for the help tho !

OpenStudy (mathstudent55):

You're welcome